# Massive tree-level splitting functions beyond kinematical limits

### v1

@@ Abstract
We present a compact form of the massive $1\to 3$ tree-level QCD splitting functions, enabling fast evaluation and improved numerical stability in practical applications. We discuss a decomposition of the results in terms of lower-order splitting functions, scalar dipole antenna functions and pure splitting remainders. The two-gluon radiator functions introduced in this context generalize the expressions obtained from the double-soft approximation.
@@ I Introduction
Final states with large multiplicities of jets initiated by massive partons are of particular interest to the future of the Large Hadron Collider (LHC) experimental programme ATLAS and CMS Collaborations (2025). Both the top quark and Higgs boson preferentially decay via bottom quarks. Higgs decays to charm quark pairs have also grown in their relevance at the high-luminosity LHC (HL-LHC), following the rapid development of jet flavor tagging algorithms powered by sophisticated machine learning techniques Qu et al. (2022); ATLAS Collaboration (2025a). Early versions of these heavy quark taggers primarily relied on identifying the decay products of heavy hadrons produced in the jets, while modern taggers exploit differences between the fragmentation patterns of light and heavy partons ATLAS Collaboration (2025b); CMS Collaboration (2024a, 2022a); ATLAS Collaboration (2025a). These algorithms have advanced to the point that processes originally thought impossible to study at the HL-LHC – namely, the production of Higgs boson pairs decaying via bottom quarks – will now be observed with high significance before the end of Run 5 ATLAS Collaboration (2025c). The development of improved theory predictions that include mass effects is thus highly motivated. At the same time, QCD jet evolution at the LHC has been measured to astonishing precision ATLAS Collaboration (2018); CMS Collaboration (2018); CMS Collaboration (2022b); ATLAS Collaboration (2020a); CMS Collaboration (2024b); ATLAS Collaboration (2024a); ALICE Collaboration (2022); ATLAS Collaboration (2025d); ALICE Collaboration (2024); CMS Collaboration (2024c); ATLAS Collaboration (2020b); ALICE Collaboration (2025). While the large dynamic range originating in the high center-of-mass energy of the collider makes it possible to treat the effects of heavy quark masses as a somewhat secondary problem today, the ever increasing experimental precision will soon lead to scenarios where massless approximations can no longer be justified. Improved theory for heavy quarks should also be included consistently within the parton shower Monte Carlo algorithms Buckley et al. (2011); Campbell et al. (2024) used both to measure jet dynamics and to train the fully supervised algorithms that empower physics analyses. This will be necessary in particular to avoid potential biases and uncertainty due to the model-dependence of classifiers ATLAS Collaboration (2024b); Gambhir et al. (2025); Garcia Caffaro et al. (2025). In this work, we compute some important components needed for precision theory in heavy quark final states: The tree-level three-parton splitting functions and the scalar dipole antenna functions including mass effects. While the splitting functions are already known Dhani et al. (2023); Craft et al. (2024), we present them in a significantly more compact form than existing literature, which will alleviate their integration into massive infrared subtraction schemes, higher logarithmic resummation and parton shower simulations, and provide faster evaluation and improved numerical stability in practical applications. We also derive, for the first time, a decomposition into scalar dipole radiators and pure splitting remainders.
@@
The factorization of QCD scattering amplitudes is at the heart of most modern techniques to compute higher-order QCD corrections to scattering processes at colliders. At fixed order in perturbation theory, it allows to construct infrared subtraction algorithms that enable the extraction of singularities from the real and virtual corrections in order to make each component finite and amenable to Monte-Carlo integration over the final-state phase space Heinrich (2021); Huss et al. (2023). In resummed perturbation theory, it enables the derivation of anomalous dimensions that govern the scaling of cross sections with changing resolution scale. For particles with masses, the factorization typically relies on the quasi-collinear limit, which requires a simultaneous scaling of the collinear transverse momentum and the parton masses. This procedure retains the correct soft behavior of collinear scattering amplitudes at leading power, but it is somewhat unnatural, because the on-shell parton masses should remain, in fact, constant. As a direct consequence, one observes a non-commutativity of limits Gaggero et al. (2022); Ghira et al. (2023, 2025) that can only be resolved by accounting for the physical features of the scattering matrix elements lost in the limiting procedure. Here we use a different concept, which avoids the introduction of kinematical limits altogether. It is based on the observation that infrared singularities arise as a consequence of degenerate asymptotic states Coleman and Norton (1965); Sterman (1995), such that QCD splitting functions can be computed as reduced matrix elements in a physical gauge Frenkel and Taylor (1976); Dokshitzer et al. (1980); Ellis et al. (1978, 1979); Humpert and van Neerven (1981). Their leading singularities are determined by the semi-classical approximation Gell-Mann and Goldberger (1954); Brown and Goble (1968), which can be derived from splitting functions in a scalar QCD, with the scalars playing the role of the semi-classical particles. This eventually allows one to derive any splitting function in terms of scalar multipole radiator functions, and pure splitting remainders with a lower degree of divergence in the kinematical limits Campbell et al. (2025).
@@ II Background, techniques and leading-order results
The motivation for our work stems from the simple yet important observation that the soft and collinear limits of QCD scattering amplitudes do not commute at leading power. To exemplify this, we consider the one-to-two parton splitting for a massive quark in the quasi-collinear limit. The splitting function reads
@@
$\langle P_{q\to q}(z)\rangle\overset{z\to 1}{\longrightarrow}C_{F}\bigg[\frac{2}{1-z}-\frac{m^{2}}{p_{q}p_{g}}\bigg]\;.$
@@
${\bf J}^{\mu}{\bf J}_{\mu}\overset{q\parallel g}{\longrightarrow}\,\propto C_{F}\bigg[\frac{2z}{1-z}-\frac{m^{2}}{p_{q}p_{g}}\bigg]\;.$
@@
The key observation of our method is that the leading components of splitting functions in a gauge theory with spin are given in terms of the semi-classical expressions, which coincide with the results of the corresponding scalar theory Gell-Mann and Goldberger (1954); Brown and Goble (1968). This allows to separate out the spin-dependent components, which do not contribute to a dipole radiation pattern. The physical picture is consistent with the well-known decomposition of the current into a charge current and a polarization current Gordon (1928). In Ref. Campbell et al. (2025), this idea was used to define a new technique to assemble splitting functions from scalar dipole radiators, lower-order splitting functions, and pure higher-order splitting remainders. Here we will extend this method to massive quarks. Because no collinear or soft limits are taken, there is no question as to how to define their extension to the quasi-collinear region, where the on-shell particle masses would have to be scaled proportional to the transverse momentum. This in turn resolves common problems which arise in the definition of polarizations of the external particles and makes the splitting functions unambiguous.
It is important to note that the splitting remainders we compute are finite in all hard collinear regions for fixed parton masses due to the dead-cone effect Marchesini and Webber (1990); Dokshitzer et al. (1991); Ghira et al. (2023). Therefore an infrared subtraction scheme can be devised without explicit knowledge of those quantities. However, for small parton masses, QCD scattering matrix elements containing such splittings develop logarithms of the hard scale over the parton masses, which should be resummed to correctly account for heavy jet dynamics. Treating such logarithms on the same footing with standard collinear poles also improves the behavior of (semi-)numerical fixed-order calculations Catani et al. (2002).
@@ II.1 Tree-level techniques
In this section we introduce the techniques needed to perform the calculation of the tree-level splitting functions. We also derive the lowest-order results, which will be used in Sec. III to decompose the splitting functions at higher order. To achieve a meaningful factorization of the results, we use physical polarization states Frenkel and Taylor (1976); Dokshitzer et al. (1980); Amati et al. (1978a, b); Ellis et al. (1978, 1979); Kalinowski et al. (1981a, b); Humpert and van Neerven (1981); Catani and Grazzini (2000). The polarization sum for virtual gluons is defined in terms of a light-like axial gauge, which is known to be ghost free Arnowitt and Fickler (1962); Konetschny and Kummer (1975, 1976); Frenkel (1976); Konetschny (1980)
@@ V.2 Gluon emission off quarks
$\begin{split}P^{({\rm pnab},p)}_{\tilde{q}\to gg\tilde{q}}&(p_{1},p_{2},p_{3})=-\frac{C_{F}C_{A}}{4}\tilde{s}_{123}^{2}\Big[\,\mathcal{S}^{\rm(nab)}_{3;3}(p_{1},p_{2};\bar{n})+\mathcal{S}^{\rm(ab)}_{3;3}(p_{1},p_{2})+2\mathcal{S}^{\rm(ab)}_{3,3;3}(p_{1},p_{2};\bar{n})\\ &\quad\;+\mathcal{S}^{\rm(ab)}_{3,3;3,3}(p_{1},p_{2};\bar{n})+\bar{\mathcal{S}}_{3;3}(p_{12};\bar{n})\Big(\mathcal{S}_{2;2}(p_{1};\bar{n})+\mathcal{S}_{1;1}(p_{2};\bar{n})-\mathcal{S}_{2;3}(p_{1};\bar{n})-\mathcal{S}_{1;3}(p_{2};\bar{n})\Big)\\ &\quad\;+2(1-\varepsilon)\,\mathcal{S}^{\mu\nu}_{3;3}(p_{12};\bar{n})\,D_{\mu}(p_{1},p_{2},\bar{n})D_{\nu}(p_{1},p_{2},\bar{n})\,\Big]+\frac{C_{A}}{2C_{F}}P^{({\rm ab},p)}_{\tilde{q}\to gg\tilde{q}}(p_{1},p_{2},p_{3})\\ &\qquad\qquad\;=C_{F}C_{A}\bigg\{\frac{\tilde{s}_{123}}{s_{12}}\bigg[\frac{\tilde{s}_{123}}{\tilde{s}_{13}}\bigg(\frac{z_{3}}{z_{2}}+\frac{z_{1}-z_{3}}{1-z_{3}}\bigg)-\frac{1-z_{1}}{z_{2}}\bigg]\\ &\;\quad-\frac{\tilde{s}_{123}}{\tilde{s}_{13}}\bigg[\frac{z_{3}(1-z_{2})}{z_{1}z_{2}}+\frac{1-z_{2}}{1-z_{3}}\bigg(1-\frac{z_{3}}{z_{1}}\bigg)-\frac{2z_{3}}{(1-z_{3})^{2}}+\frac{2m^{2}}{s_{12}}\frac{z_{1}^{2}}{z_{2}(1-z_{3})}\bigg]-\frac{s_{12}}{\tilde{s}_{13}}\frac{2z_{3}(1-z_{2})}{z_{1}(1-z_{3})^{2}}\\ &\;\quad+\frac{2z_{3}(z_{1}-z_{2})}{z_{2}(1-z_{3})^{2}}+\frac{2z_{3}}{z_{1}z_{2}}+\frac{1-\varepsilon}{4}+\frac{m^{2}}{s_{12}}\bigg(1+\frac{2z_{1}^{2}}{z_{2}(1-z_{3})}\bigg)-\frac{2m^{2}}{\tilde{s}_{13}}\frac{1}{z_{2}}\bigg\}+(1\leftrightarrow 2)\;.\end{split}$ (50)
@@
$\begin{split}&\langle P^{({\rm pnab},p,{\rm f})}_{q\to ggq}(p_{1},p_{2},p_{3})\rangle=C_{F}C_{A}\,(1-\varepsilon)\,\bigg\{\frac{z_{2}}{z_{1}}\bigg(\frac{1}{1-z_{1}}+\frac{z_{3}}{1-z_{2}}\bigg)+\bigg(\frac{2z_{2}}{z_{1}}+1\bigg)\bigg(1-\frac{3}{4}\frac{\tilde{s}_{123}}{s_{12}}(1-z_{3})\bigg)\\ &\;\qquad+\frac{(1-\varepsilon)}{2}\bigg[\,1+\frac{s_{12}}{\tilde{s}_{13}}+\frac{z_{1}}{1-z_{2}}\bigg(\frac{z_{2}}{1-z_{2}}-\frac{\tilde{s}_{123}}{\tilde{s}_{13}}\bigg)\bigg]+\frac{\tilde{s}_{123}}{2\tilde{s}_{13}}\bigg(\frac{\tilde{s}_{123}}{s_{12}}-\frac{1-z_{2}}{z_{1}}\bigg)\bigg(\frac{(1-z_{3})^{2}}{z_{2}}+\frac{z_{2}^{2}}{1-z_{3}}\bigg)\\ &\;\qquad+\bigg(\frac{\tilde{s}_{123}}{s_{12}}\frac{1-z_{3}}{z_{1}}+\frac{\tilde{s}_{123}}{\tilde{s}_{13}}\frac{1-z_{2}}{z_{1}}-\frac{\tilde{s}_{123}^{2}}{s_{12}\tilde{s}_{13}}\bigg)\bigg(1-z_{3}-\frac{s_{12}}{\tilde{s}_{123}}\bigg)+\frac{m^{2}}{\tilde{s}_{13}}\frac{z_{1}-z_{2}}{1-z_{2}}-\frac{3}{4}\bigg\}+(1\leftrightarrow 2)\;.\end{split}$ (51)

### v2

@@ Abstract
We present a compact form of the massive $1\to 3$ tree-level QCD splitting functions and discuss a decomposition of the results in terms of lower-order expressions, scalar dipole antenna functions and pure higher-order remainders. The two-gluon radiator functions introduced in this context are novel and generalize expressions obtained from the double-soft approximation. Our results are obtained without reference to soft or quasi-collinear limits.
@@ I Introduction
Final states with large multiplicities of jets initiated by massive partons are of particular interest to the future of the Large Hadron Collider (LHC) experimental programme ATLAS and CMS Collaborations (2025). Both the top quark and Higgs boson preferentially decay via bottom quarks. Higgs decays to charm quark pairs have also grown in their relevance at the high-luminosity LHC (HL-LHC), following the rapid development of jet flavor tagging algorithms powered by sophisticated machine learning techniques Qu et al. (2022); ATLAS Collaboration (2025a). Early versions of these heavy quark taggers primarily relied on identifying the decay products of heavy hadrons produced in the jets, while modern taggers exploit differences between the fragmentation patterns of light and heavy partons ATLAS Collaboration (2025b); CMS Collaboration (2024a, 2022a); ATLAS Collaboration (2025a). These algorithms have advanced to the point that processes originally thought impossible to study at the HL-LHC – namely, the production of Higgs boson pairs decaying via bottom quarks – will now be observed with high significance before the end of Run 5 ATLAS Collaboration (2025c). The development of improved theory predictions that include mass effects is thus highly motivated. At the same time, QCD jet evolution at the LHC has been measured to astonishing precision ATLAS Collaboration (2018); CMS Collaboration (2018); CMS Collaboration (2022b); ATLAS Collaboration (2020a); CMS Collaboration (2024b); ATLAS Collaboration (2024a); ALICE Collaboration (2022); ATLAS Collaboration (2025d); ALICE Collaboration (2024); CMS Collaboration (2024c); ATLAS Collaboration (2020b); ALICE Collaboration (2025). While the large dynamic range originating in the high center-of-mass energy of the collider makes it possible to treat the effects of heavy quark masses as a somewhat secondary problem today, the ever increasing experimental precision will soon lead to scenarios where massless approximations can no longer be justified. Improved theory for heavy quarks should also be included consistently within the parton shower Monte Carlo algorithms Buckley et al. (2011); Campbell et al. (2024) used both to measure jet dynamics and to train the fully supervised algorithms that empower physics analyses. This will be necessary in particular to avoid potential biases and uncertainty due to the model-dependence of classifiers ATLAS Collaboration (2024b); Gambhir et al. (2025); Garcia Caffaro et al. (2025). In this work, we compute some important components needed for precision theory in heavy quark final states: The tree-level three-parton splitting functions and the scalar dipole antenna functions including mass effects. While the splitting functions are already known Dhani et al. (2023); Craft et al. (2024), we present them in a significantly more compact form than existing literature, which will alleviate their integration into massive infrared subtraction schemes, higher logarithmic resummation and parton shower simulations, and provide faster evaluation and improved numerical stability in practical applications. The massive scalar dipole radiators are computed for the first time, as is the decomposition of the splitting functions into scalar radiators and pure splitting remainders.
@@
The factorization of QCD scattering amplitudes is at the heart of most modern techniques to compute higher-order QCD corrections to scattering processes at colliders. At fixed order in perturbation theory, it allows one to construct infrared subtraction algorithms that enable the extraction of singularities from the real and virtual corrections in order to make each component finite and amenable to Monte-Carlo integration over the final-state phase space Heinrich (2021); Huss et al. (2023). In resummed perturbation theory, it enables the derivation of anomalous dimensions that govern the scaling of cross sections with changing resolution scale. For particles with masses, the factorization typically relies on the quasi-collinear limit, which requires a simultaneous scaling of the collinear transverse momentum and the parton masses. This procedure retains the correct soft behavior of collinear scattering amplitudes at leading power, but it is somewhat unphysical, because the on-shell parton masses should remain, in fact, constant. As a direct consequence, one observes a non-commutativity of limits Gaggero et al. (2022); Ghira et al. (2023, 2025) that can only be resolved by accounting for the physical features of the scattering matrix elements lost in the limiting procedure. Here we use a different concept, which avoids the introduction of kinematical limits altogether. It is based on the observation that infrared singularities arise as a consequence of degenerate asymptotic states Coleman and Norton (1965); Sterman (1995), such that QCD splitting functions can be computed as reduced matrix elements in a physical gauge Frenkel and Taylor (1976); Dokshitzer et al. (1980); Ellis et al. (1978, 1979); Humpert and van Neerven (1981). Their leading singularities are determined by the semi-classical approximation Gell-Mann and Goldberger (1954); Brown and Goble (1968), which can be derived from splitting functions in a scalar QCD, with the scalars playing the role of the semi-classical particles. This eventually allows one to derive any splitting function in terms of scalar multipole radiator functions, and pure splitting remainders with a lower degree of divergence in the kinematical limits Campbell et al. (2025).
@@ II Background, techniques and leading-order results
The motivation for our work stems from the simple yet important observation that the soft and collinear limits of QCD scattering amplitudes do not commute beyond leading power. To exemplify this, we consider the one-to-two parton splitting for a massive quark in the quasi-collinear limit. The splitting function reads
@@
$\langle P_{q\to q}(z)\rangle\overset{z\to 1}{\longrightarrow}C_{F}\bigg[\frac{2}{1-z}-\frac{m_{q}^{2}}{p_{q}p_{g}}\bigg]\;.$
@@
${\bf J}^{\mu}{\bf J}_{\mu}\overset{q\parallel g}{\longrightarrow}\,\propto C_{F}\bigg[\frac{2z}{1-z}-\frac{m_{q}^{2}}{p_{q}p_{g}}\bigg]\;.$
@@
The key observation of our method is that the leading components of splitting functions in a gauge theory with spin are given in terms of the semi-classical expressions, which coincide with the results of the corresponding scalar theory Gell-Mann and Goldberger (1954); Brown and Goble (1968). This allows us to separate out the spin-dependent components, which do not contribute to a dipole radiation pattern. The physical picture is consistent with the well-known decomposition of the current into a charge current and a polarization current Gordon (1928). In Ref. Campbell et al. (2025), this idea was used to define a new technique to assemble splitting functions from scalar dipole radiators, lower-order splitting functions, and pure higher-order splitting remainders. Here we will extend this method to massive quarks. Because no collinear or soft limits are taken, there is no question as to how to define their extension to the quasi-collinear region, where the on-shell particle masses would have to be scaled proportional to the transverse momentum. This in turn resolves common problems which arise in the definition of polarizations of the external particles and makes the splitting functions unambiguous.
It is important to note that the splitting remainders we compute are finite in all hard collinear regions for fixed parton masses due to the dead-cone effect Marchesini and Webber (1990); Dokshitzer et al. (1991); Ghira et al. (2023). Therefore, an infrared subtraction scheme can be devised without explicit knowledge of those quantities. However, for small parton masses, QCD scattering matrix elements containing such splittings develop logarithms of the hard scale over the parton masses, which should be resummed to correctly account for heavy jet dynamics. Treating such logarithms on the same footing with standard collinear poles also improves the behavior of (semi-)numerical fixed-order calculations Catani et al. (2002).
@@ II.1 Tree-level techniques
In this section we introduce the techniques needed to perform the calculation of the tree-level splitting functions. We also derive the lowest-order results, which will be used in Sec. III to decompose the splitting functions at higher order. Similar approaches, working directly at the level of helicity amplitudes, were introduced in Duhr and Maltoni (2008); Cohen and Riembau (2024). To achieve a meaningful factorization of the results, we use physical polarization states Frenkel and Taylor (1976); Dokshitzer et al. (1980); Amati et al. (1978a, b); Ellis et al. (1978, 1979); Kalinowski et al. (1981a, b); Humpert and van Neerven (1981); Catani and Grazzini (2000). The polarization sum for virtual gluons is defined in terms of a light-like axial gauge, which is known to be ghost free Arnowitt and Fickler (1962); Konetschny and Kummer (1975, 1976); Frenkel (1976); Konetschny (1980)
@@ V.2 Gluon emission off quarks
$\begin{split}P^{({\rm pnab},p)}_{\tilde{q}\to gg\tilde{q}}&(p_{1},p_{2},p_{3})=-\frac{C_{F}C_{A}}{4}\tilde{s}_{123}^{2}\Big[\,\mathcal{S}^{\rm(nab)}_{3;3}(p_{1},p_{2};\bar{n})+\mathcal{S}^{\rm(ab)}_{3;3}(p_{1},p_{2})+2\mathcal{S}^{\rm(ab)}_{3,3;3}(p_{1},p_{2};\bar{n})\\ &\quad\;+\mathcal{S}^{\rm(ab)}_{3,3;3,3}(p_{1},p_{2};\bar{n})+\bar{\mathcal{S}}_{3;3}(p_{12};\bar{n})\Big(\mathcal{S}_{2;2}(p_{1};\bar{n})+\mathcal{S}_{1;1}(p_{2};\bar{n})-\mathcal{S}_{2;3}(p_{1};\bar{n})-\mathcal{S}_{1;3}(p_{2};\bar{n})\Big)\\ &\quad\;+2(1-\varepsilon)\,\mathcal{S}^{\mu\nu}_{3;3}(p_{12};\bar{n})\,D_{\mu}(p_{1},p_{2},\bar{n})D_{\nu}(p_{1},p_{2},\bar{n})\,\Big]+\frac{C_{A}}{2C_{F}}P^{({\rm ab},p)}_{\tilde{q}\to gg\tilde{q}}(p_{1},p_{2},p_{3})\\ &\qquad\qquad\;=C_{F}C_{A}\bigg\{\frac{1-\varepsilon}{4}+\bigg(\frac{\tilde{s}_{123}}{\tilde{s}_{13}}\frac{z_{1}}{1-z_{3}}-1\bigg)\bigg(\frac{\tilde{s}_{123}}{s_{12}}\frac{1-z_{1}}{z_{2}}-\frac{2m^{2}}{s_{12}}\frac{z_{1}}{z_{2}}-1\bigg)-1\\ &\;\quad+\bigg[\frac{\tilde{s}_{13}-\tilde{s}_{23}}{\tilde{s}_{13}}\frac{z_{1}-z_{2}}{z_{2}}-\frac{s_{12}}{\tilde{s}_{13}}\bigg(\frac{1-z_{3}}{z_{2}}+\frac{2z_{3}}{z_{1}}\bigg)\bigg]\frac{z_{3}}{(1-z_{3})^{2}}+\frac{2z_{3}}{z_{1}z_{2}}-\frac{2m^{2}}{\tilde{s}_{13}}\frac{1}{z_{2}}\bigg\}+(1\leftrightarrow 2)\;.\end{split}$ (50)
@@
$\begin{split}&\langle P^{({\rm pnab},p,{\rm f})}_{q\to ggq}(p_{1},p_{2},p_{3})\rangle=C_{F}C_{A}\,(1-\varepsilon)\,\bigg\{\frac{z_{2}}{z_{1}}\bigg(\frac{1}{1-z_{1}}+\frac{z_{3}}{1-z_{2}}\bigg)+\bigg(\frac{2z_{2}}{z_{1}}+1\bigg)\bigg(1-\frac{3}{4}\frac{\tilde{s}_{123}}{s_{12}}(1-z_{3})\bigg)\\ &\;\qquad+\frac{1-\varepsilon}{2}\bigg[\,1+\frac{s_{12}}{\tilde{s}_{13}}+\frac{z_{1}}{1-z_{2}}\bigg(\frac{z_{2}}{1-z_{2}}-\frac{\tilde{s}_{123}}{\tilde{s}_{13}}\bigg)\bigg]+\frac{\tilde{s}_{123}}{2\tilde{s}_{13}}\bigg(\frac{\tilde{s}_{123}}{s_{12}}-\frac{1-z_{2}}{z_{1}}\bigg)\bigg(\frac{(1-z_{3})^{2}}{z_{2}}+\frac{z_{2}^{2}}{1-z_{3}}\bigg)\\ &\;\qquad+\bigg(\frac{\tilde{s}_{123}}{s_{12}}\frac{1-z_{3}}{z_{1}}+\frac{\tilde{s}_{123}}{\tilde{s}_{13}}\frac{1-z_{2}}{z_{1}}-\frac{\tilde{s}_{123}^{2}}{s_{12}\tilde{s}_{13}}\bigg)\bigg(1-z_{3}-\frac{s_{12}}{\tilde{s}_{123}}\bigg)+\frac{m^{2}}{\tilde{s}_{13}}\frac{z_{1}-z_{2}}{1-z_{2}}-\frac{3}{4}\bigg\}+(1\leftrightarrow 2)\;.\end{split}$ (51)